# United Minds or Isolated Agents? Exploring Coordination of LLMs under Cognitive Load Theory

### v1

@@ 5.4 Ablation Studies on LiveBench Subsets
Figure 4 demonstrates CoThinker’s hyperparameter sensitivity, offering insights into cognitive load management as theorized in Section 3. The reference set size ($N$, top row) directly impacts extraneous cognitive load. An optimal $N$ (e.g., 2-3) balances diverse peer input against cognitive overload, respecting LLM working memory limits. Too few references limit collaboration; too many overwhelm. The exploration rate ($\beta$, middle row) governs the trade-off between exploiting similar ideas (low $\beta$, lower extraneous load for integration) and exploring diverse ones (high $\beta$, high extraneous load). Task-dependent optima, like higher $\beta$ for Reasoning, reflect this balance, managed by the Communication Moderator’s cognitive small-world network. The number of agents ($M$, bottom row) shows that while more agents can distribute intrinsic load, increasing $M$ also elevates transactional (extraneous) load from coordination. Non-monotonic performance indicates that beyond a point, these transactional costs negate the benefits of parallelism, aligning with CLT’s predictions for group overload. These findings affirm that CoThinker’s parameters are crucial for managing cognitive load, enabling the emergence of an effective "collective mind" by mitigating overload.
@@ 6 Conclusion
This work addresses the performance limitations of LLMs on complex tasks, particularly when employing In-Context Learning (ICL), by drawing an analogy to Cognitive Load Theory (CLT). We posit that observed performance ceilings arise from exceeding an LLM’s effective cognitive load capacity when processing intricate task details and extensive in-context guidance. We introduced CoThinker, a multi-agent architecture that operationalizes CLT principles. Through agent specialization, a transactive memory system, and moderated communication, CoThinker mitigates overload and enhances collaborative problem-solving, especially for tasks that challenge single agents using ICL. Empirical evaluations on benchmarks like LiveBench and CommonGen-Hard demonstrated CoThinker’s superior performance over existing baselines on high-load tasks. Analyses validated CoThinker’s effective management of cognitive load, fostering a more robust "collective mind." By grounding multi-agent LLM design in CLT, this research offers a principled path towards overcoming performance bottlenecks encountered when applying ICL to demanding problems, contributing to more powerful collaborative AI systems through the lens of cognitive science.
@@
Appendix A Cognitive Foundations: Elaborations
A.1 Human Working Memory and Attentional Control
@@
A.2 Using Cognitive Load Theory to Explain Phenomena in LLM Performance
Cognitive Load Theory (CLT) offers a valuable lens to interpret puzzling LLM performance issues, positing that LLMs, like humans, have finite processing capacity. Exceeding this capacity leads to performance degradation. This section concisely analyzes several such cases through CLT.
@@ A.2 Using Cognitive Load Theory to Explain Phenomena in LLM Performance
Appendix B CoThinker Architecture: Implementation and Prompting
B.1 Prompt Architecture for Agent Parallel Thinking
@@
B.2 Prompt Architecture for Transactive Memory System (TMS) Emulation
As introduced in Section 4.2, CoThinker incorporates a mechanism to emulate a human Transactive Memory System (TMS). A TMS is a collective cognitive resource developed by groups, encompassing a shared understanding of who knows what (metamemory or expertise directory), how to access and integrate this distributed knowledge, and a level of trust in the information provided by different members (Wegner, 1987; Hollingshead, 2001; Lewis, 2003). Effective TMS functioning involves processes of knowledge encoding (assigning information to members or recognizing expertise), storage (individuals retaining specialized knowledge), and retrieval (accessing and using the distributed knowledge), facilitated by member specialization, perceived credibility, and inter-agent coordination (Yoo and Kanawattanachai, 2001). This systematic division and integration of cognitive labor allows groups to handle more complex information and solve problems more effectively than individuals or less coordinated groups.
@@ B.2 Prompt Architecture for Transactive Memory System (TMS) Emulation
B.3 Communication Moderator: Cultivating an Efficient Network via Strong and Weak Ties
The Communication Moderator in CoThinker (Section 4.3) strategically structures inter-agent communication by implicitly leveraging principles from social and complex network theories. This design fosters a network optimized for managing cognitive load and enhancing collective intelligence.
@@ B.3 Communication Moderator: Cultivating an Efficient Network via Strong and Weak Ties
  • Network Outcome: Introducing such weak ties into a highly clustered network is a hallmark of small-world networks (Watts and Strogatz, 1998). These networks advantageously combine high local clustering with short global average path lengths.
@@
B.4 Synthesizer Module: Consolidation and Cognitive Grounding
The Synthesizer module (Section 4.4) consolidates outputs from all agents ($\{x_{i}^{(T-1)}\}_{i=1}^{M}$) and the final Transactive Memory System state ($\mu^{(T-1)}$) into a single solution for the task $D$. The design choice for the Synthesizer can vary, with different cognitive implications:

### v2

@@ 5.5 CoThinker Ablation of Communication Moderator
We ablate the Communication Moderator’s key parameters—reference set size ($N$), exploration rate ($\beta$), and agent count (MM)—on Gemini-1.5-Flash-8B across four LiveBench categories: Math, Reasoning, Data Analysis, and Instruction (See Figure 4). Default parameters: $T=3$, with controlled variations for each parameter. Scores are normalized by IO baseline. Analysis. The Communication Moderator’s parameters directly control CL balance: Reference set size ($N$) manages extraneous load—optimal N=23N=2{-}3 balances peer input diversity against overload, respecting LLM working memory limits. Exploration rate ($\beta$) governs similarity-diversity trade-offs: low $\beta$ (exploiting similar ideas) reduces integration load but risks echo chambers; high $\beta$ (exploring diverse perspectives) aids intrinsic load distribution but increases extraneous load. Task-dependent optima (e.g., higher $\beta$ for Reasoning) reflect this balance through small-world network properties. Agent count ($M$) shows non-monotonic performance—more agents distribute intrinsic load but elevate coordination costs, confirming CLT predictions for group overload. These findings validate the Communication Moderator’s role in managing CL for effective collective intelligence. See Appendix F for details.
Figure 4: Communication Moderator ablation on CoThinker using Gemini-1.5-Flash-8B. Top: Reference Set Size ($N\in\{0,2,3,4,5\}$, $M=6,\beta=0.3$); Middle: Exploration Rate ($\beta\in\{0.1,0.3,0.6,1.0\}$, $N=2,M=6$); Bottom: Agent Count ($M\in\{6,12,18\}$, $N=3,\beta=0.3$).
5.6 CoThinker Ablation on Other Components
We conducted component ablation on Transactive Memory System (TMS) and Thinking Style Orchestrator (Style) a subset of LiveBench Math tasks. We also examine the proxy, perplexity (PPL), to reflect the CL management effects of our components.
Component Ablation. Communication Moderator is fixed ON for all runs. We test configurations: TMS $\in$ {On, Off}; Style $\in$ {On, Off}. Analysis. Our components benefit most models (details in Appendix E.4). From Table 4, we find Thinking Style Orchestrator provides consistent improvements. However, TMS is less effective for new GPT models. Investigating their output, we find that they often refuse to give intermediate results, responding with "I can’t provide step-by-step reasoning." This is counterproductive, as it discourages the detailed reasoning used to build TMS.
Configuration Qwen3-30B-A3B GPT5-Nano GPT-OSS-20B
TMS: ON, Styles: ON 81.87 55.97 57.15
TMS: ON, Styles: OFF 69.79 49.02 48.21
TMS: OFF, Styles: ON 76.41 62.36 58.37
Table 4: Component ablation on subset of Math dataset, with effect of TMS and Thinking Style Orchestrator, indicating the CL management benefits of each component.
PPL proxy as Evidence.. We conducted perplexity (PPL) studies on weaker models to demonstrate how our components helps weaker models reduce CL in understanding stronger models’ outputs. We choose Math and Reasoning tasks for our analysis. Lower PPL indicates higher easiness and effective CL reduction (See Table 5). For more interesting ablation with PPL proxy, see Appendix E.5.
Model Baseline Styles TMS References (N=3)
Qwen3-8B 6.56 3.58 1.69 3.10
Mistral-7B 6.63 5.04 1.58 2.86
Table 5: PPL ablation showing CL reduction effects: how components help weaker models process information from better models’ answers, reducing CL (lower PPL).
@@ 6 Conclusion
This work explains ICL performance ceilings in LLMs through Cognitive Load Theory (CLT), attributing failures on complex tasks to cognitive overload. We establish both theoretical foundations mapping human working memory to LLM attention mechanisms and empirical examination through measurable cognitive load proxies (attention entropy and perplexity), demonstrating that the CLT-LLM analogy provides a principled framework for understanding and addressing performance limitations. We introduce CoThinker, a CLT-grounded multi-agent architecture with specialized agents, a transactive memory, and communication moderator to reduce load. On benchmarks with complex tasks, CoThinker consistently outperforms strong baselines. Analysis confirm effective cognitive load management, suggesting a principled CLT-based path to more capable collaborative LLM systems.
Ethics Statement
This research contributes to society and human well-being by advancing scientific understanding of cognitive load principles in AI systems, potentially enabling more capable and efficient collaborative AI. We uphold high standards of scientific excellence through rigorous experimental design, transparent reporting of methods and results, and honest acknowledgment of limitations. The work avoids harm by focusing on computational improvements without involving human subjects, sensitive personal data, or applications with direct societal risks. We maintain honesty and transparency by providing complete implementation details, acknowledging all limitations, and declaring no conflicts of interest. The research promotes fairness by using publicly available benchmarks and established evaluation protocols accessible to the broader research community. We respect intellectual property by properly citing all prior work and acknowledging the foundation provided by existing research. Privacy is honored as no personal data collection or processing occurs. All experimental data uses publicly available benchmarks and API-based LLM services under their respective terms of service.
Reproducibility Statement
Our methods rely primarily on existing LLM APIs where we have detailed the specific models used and provided sample prompts and exact agent workflow descriptions in the main paper and appendix. Complete implementation details for reproducing our results are provided, including all agent prompts, communication protocols, system configurations, and evaluation procedures. The benchmark datasets (LiveBench, CommonGen-Hard) are publicly available with standardized evaluation metrics that enable direct comparison.
@@
Appendix A LLM Usage
LLMs served as experimental subjects for evaluating our cognitive load theory hypotheses (not as sources of research ideation). They were additionally used in limited code authoring support and writing assistance (grammar correction and standard-English rephrasing). They were not used for generating core methodological ideas, designing experiments, forming hypotheses, performing analysis, or implementing core architectural components. All conceptual framing, system design, analysis, and conclusions are entirely human-generated. Authors retain full responsibility for all content.
Appendix B Cognitive Foundations and Analogy Elaboration
Here we elaborate on the cognitive science foundations underpinning our framework and the analogy to LLMs from Section 3.1.
B.1 Human Working Memory and Attentional Control
@@
B.2 Working Memory and Parametric Knowledge in LLMs
Misunderstanding 1: Working Memory as Context Window Limits A prevalent misunderstanding in current LLM research equates Working Memory (WM) with physical limitations like context window size and assumes that models only fail when they exceed their maximum token limits—that WM is simply the amount of in-context information that can fit within the context window. However, our framework distinguishes WM as the mechanism of selective attention: a limited cognitive capacity to actively hold and manipulate information simultaneously, not merely passive storage of tokens. This selective attention mechanism can become overwhelmed and lead to performance degradation well before reaching physical context limits.
Misunderstanding 2: High Cognitive Load as Long Input/Output Tasks Another common misconception is that high cognitive load (CL) tasks are simply those with long inputs and long outputs. This conflates task length with cognitive complexity. In our framework, a high CL task is defined by the working memory demands it places on simultaneous information processing—specifically, the amount of information that must be actively held and manipulated simultaneously in WM. For example, a sequential fact elicitation task (e.g., "List 100 countries and their capitals") may have long input and output but requires minimal WM since each fact can be retrieved independently without integrating interdependent information. Conversely, a complex reasoning task requiring simultaneous consideration of multiple interacting constraints, relationships, or variables creates high CL regardless of input/output length, as it demands that the model maintain and manipulate multiple information pieces concurrently in its working memory.
Our Framework: Working Memory and Parametric Knowledge Our framework clarifies that an LLM’s parameters can be understood as 1)Parametric Knowledge (analogous to human long-term memory, mainly from FFN layer Geva et al. (2021)), and 2) innate generation capabilities that determine how effectively the model can select and process information for its WM (mainly attention layer Tighidet et al. (2025)). A model’s parametric knowledge affects how well it understands and integrates information during reasoning (reflected in attention patterns), while WM serves as the active processing resource that must load both contextual information and activated parametric knowledge Tighidet et al. (2025). Complex tasks demanding simultaneous consideration of multiple interacting elements can cause cognitive overload regardless of context window utilization or task length. Our pilot study (Appendix C) provides empirical validation of this WM framework through measurable cognitive load proxies.
B.3 Detailed Mechanisms of Human Collective Intelligence
Here we elaborate on the cognitive mechanisms enabling human collective intelligence from Section 3.1. Human collective intelligence emerges from sophisticated social-cognitive abilities that enable groups to surpass individual cognitive limitations through several key mechanisms:
Shared Intentionality and Theory of Mind: Effective collective intelligence requires individuals to understand others’ mental states and coordinate intentions toward common goals (Tomasello et al., 2005; Frith & Frith, 2005), enabling the establishment of common ground necessary for distributed cognitive processing.
Meta-cognitive Awareness: Individuals develop meta-knowledge about "who knows what" (Hollingshead, 2001), enabling efficient allocation of cognitive resources and allowing group members to rely on each other for information sharing and retrieval (Hollingshead & Brandon, 2003).
Spontaneous Organization: Effective collective intelligence often emerges spontaneously through self-organizing principles (Shteynberg et al., 2023), with groups naturally developing communication patterns and role distributions that optimize cognitive load management.
Structured Communication: Groups develop specialized communication protocols that establish common ground, minimizing extraneous cognitive load while maximizing information integration (Tomasello et al., 2005).
These mechanisms demonstrate that human collective intelligence results from emergent properties of social-cognitive interaction that specifically address cognitive load management challenges, providing natural solutions to cognitive overload that inform LLM multi-agent system design.
B.4 Using Cognitive Load Theory to Explain Phenomena in LLM Performance
Here we provide examples of how CLT explains LLM performance phenomena from Section 3.1. Cognitive Load Theory (CLT) offers a valuable lens to interpret puzzling LLM performance issues, positing that LLMs, like humans, have finite processing capacity. Exceeding this capacity leads to performance degradation. This section concisely analyzes several such cases through CLT.
@@ B.4 Using Cognitive Load Theory to Explain Phenomena in LLM Performance
Appendix C Pilot Study: Empirical Validation of CLT-LLM Analogy
This section provides detailed implementation and analysis of our pilot study that empirically validates the Cognitive Load Theory (CLT) analogy for LLMs, as summarized in Section 3.2.
C.1 Mathematical Definitions of Cognitive Load Proxies
Attention Entropy measures the diversity of the model’s attention distribution, formally defined as:
$H=-\sum_{i=1}^{N}a_{i}\log a_{i}$ (7)
where $a_{i}$ is the normalized attention weight on token $i$ and $N$ is the total number of tokens. Higher entropy indicates more uniform attention distribution across input tokens, suggesting the model must consider more aspects of the input simultaneously.
Perplexity measures the model’s uncertainty about its predictions, formally defined as:
$\text{PPL}=\exp\left({-\frac{1}{N}\sum_{i=1}^{N}\log P(w_{i})}\right)$ (8)
where $P(w_{i})$ is the probability assigned to token $w_{i}$ and $N$ is the sequence length. Lower perplexity indicates higher confidence in predictions.
C.2 Theoretical Justification for Proxies
The two proxies we employ actually measure LLM-perceived cognitive load that must be accommodated within the model’s working memory for task completion. This distinction is crucial for understanding what we are measuring:
Attention Entropy as Working Memory Load: By definition, working memory involves the active manipulation and maintenance of information through attentional control mechanisms. When a task requires considering multiple, disparate pieces of information simultaneously, attention must be distributed across these elements, resulting in higher entropy. This distributed attention pattern reflects the cognitive effort required to maintain multiple information pieces in working memory, directly corresponding to intrinsic cognitive load as defined by CLT.
Perplexity as Task Easiness: Cognitive load theory posits that when cognitive demands exceed working memory capacity, performance degrades and uncertainty increases. Perplexity captures this uncertainty - when a model is cognitively overloaded, it becomes less confident about its responses, manifesting as higher perplexity. Conversely, tasks within the model’s cognitive capacity yield confident, low-perplexity responses.
C.3 Experimental Setup
Model: We conducted experiments using Mistral-7B-v0.3, a mid-sized language model that exhibits clear cognitive limitations while maintaining reasonable performance across diverse tasks.
Prefilled vs. Generated Answers: We measure perplexity on prefilled ground-truth answers rather than model-generated responses to avoid heuristic confounds: cognitively overloaded models may produce confident but incorrect responses (Agarwal et al., 2024), which would not reflect the true cognitive load of the underlying task.
Implementation Details: For Attention Entropy, we aggregate softmax attention weights, then average over heads and layers.
C.4 Experiment 1: Attention Entropy and Task Complexity
Dataset Construction: We constructed a controlled dataset from AMPS-Hard, focusing on arithmetic reasoning tasks. To ensure fair comparison, we:
  • Selected 4 difficulty levels (simple arithmetic to complex multi-step problems)
  • Controlled input length across difficulty levels to isolate complexity effects
  • Ensured consistent question format and domain (mathematical reasoning)
Different question types are not directly comparable using attention entropy, as the metric is sensitive to input structure and content. Therefore, we focused on a single domain with graduated difficulty.
Task Complexity Attention Entropy (No Reasoning) Attention Entropy (With Reasoning)
Level 1 4.442 4.439
Level 2 4.796 4.726
Level 3 5.043 4.937
Level 4 6.101 5.920
Table 6: Attention entropy increases with task complexity, validating the proxy. Reasoning steps reduce entropy by helping the model focus on key information.
Analysis. Attention entropy increases monotonically with task complexity, indicating that harder tasks require the model to consider more information pieces simultaneously, corresponding to higher cognitive load. Notably, when reasoning steps are provided, attention entropy decreases across all levels, suggesting that explicit reasoning helps the model "offload" partial cognitive load by providing a structured focus on key information elements.
C.5 Experiment 2: Perplexity and Instruction Complexity
Dataset Construction: We used the FLASK dataset, specifically filtering for problems within the same category to ensure comparability. We selected pairs of problems where one was marked as requiring "expert knowledge" and another without this marking, representing hard and easy tasks respectively within the same domain.
Instruction Complexity Levels:
  • Level 1: Blank (no additional instruction)
  • Level 2: "Think step by step"
  • Level 3: "Please think step by step, focusing on factuality and logical reasoning"
  • Level 4: Extended reasoning guidance with specific cognitive strategies
  • Level 5: Comprehensive instruction with multiple reasoning frameworks
Analysis. For hard tasks, perplexity initially decreases with instruction complexity (Levels 1-3), indicating that structured guidance helps the model focus and reduces cognitive uncertainty. However, perplexity increases again at higher instruction levels (Levels 4-5), suggesting that excessive instruction becomes an additional cognitive burden - classic extraneous load as predicted by CLT.
For easy tasks, perplexity remains consistently low and slightly increases with instruction complexity, indicating that additional guidance provides no benefit and may even introduce unnecessary extraneous load. This pattern perfectly aligns with CLT predictions about the differential effects of instructional design on tasks of varying intrinsic difficulty.
Instruction Complexity Perplexity (Hard) Perplexity (Easy)
Level 1 120.50 3.37
Level 2 88.97 3.42
Level 3 85.35 3.45
Level 4 92.48 3.46
Level 5 100.71 3.46
Table 7: Perplexity patterns validate CLT predictions across instruction complexity levels.
C.6 Proxy Scope: Diagnostic, Not Operational
Why not test-time use. The proxies we study (attention entropy and perplexity) are not universally applicable at inference time. First, when cognitively overloaded, models may fall back on heuristics, yielding confident but incorrect responses (Agarwal et al., 2024); this breaks the assumption that a model’s own outputs reveal the true load of the underlying task. Second, perplexity requires ground-truth answers, which are unavailable during inference. Together, these factors also clarify why using PPL as a test-time assistant tends to fail: heuristic-driven outputs can appear easy for the model while still being wrong, and ground-truth-based PPL cannot be computed online.
How we ensured clean experimental signals. In our experiments, we treat these proxies strictly as diagnostic rather than operational signals. To mitigate confounds, we (i) controlled task domain and type within each study (addressing cross-task incomparability), (ii) controlled input length within each complexity level (addressing length effects on both metrics), and (iii) evaluated on prefilled ground-truth answers rather than model-generated outputs to avoid heuristic confounds. Consequently, the proxies serve to validate the theoretical foundations, not to drive the system at test time—hence CoThinker emphasizes architectural solutions over real-time cognitive load monitoring.
Appendix D CoThinker Architecture: Implementation and Prompting
D.1 CoThinker as Plug-and-Play CLT Principles for Multi-Agent Frameworks
CoThinker’s primary contribution lies in providing cognitive design principles derived from Cognitive Load Theory (CLT) that can be integrated into existing multi-agent frameworks as "plug-and-play" enhancements. Rather than competing with established platforms, CoThinker offers a complementary cognitive lens that addresses a fundamental gap: most existing frameworks focus on engineering coordination without explicitly managing cognitive load.
The multi-agent LLM ecosystem encompasses diverse frameworks serving different purposes: Engineering Platforms like AutoGen (Wu et al., 2023) and CrewAI focus on conversational flows and role-based task delegation; Domain-Specific Systems like MetaGPT (Hong et al., 2023) and ChatDev (Qian et al., 2023b) target software development workflows; Research Frameworks like AgentVerse (Chen et al., 2023b) enable experimental exploration; Workflow Systems like LangGraph provide stateful execution; and Consensus Systems like Multi-Agent Debate (Du et al., 2023) and ReConcile (Chen et al., 2023a) focus on group reasoning.
Unlike these frameworks that primarily address engineering coordination, CoThinker introduces explicit cognitive load management through CLT-grounded design principles. Our approach provides three core cognitive enhancements: (1) dynamic thinking style assignment based on task demands rather than fixed functional roles, (2) communication moderation that balances cognitive similarity and diversity to prevent overload, and (3) transactive memory systems that enable cognitive offloading and specialization. These principles can augment existing frameworks—for instance, enhancing AutoGen’s group chat with load-aware communication, replacing CrewAI’s static roles with adaptive cognitive styles, or integrating TMS-based knowledge sharing into MetaGPT’s development workflows.
D.2 Information Flow and Message Passing
This section provides detailed clarification of how information is generated and passed through the CoThinker system, addressing reviewer questions about system dynamics.
Multi-Channel Information Architecture: Each agent receives input through two distinct channels designed to manage cognitive load:
  • TMS Channel: Provides high-level, synthesized, long-term context (group progress, consensus, "who knows what"), giving agents strategic awareness with low cognitive burden.
  • Communication Moderator Channel: Provides specific, low-volume, current peer outputs for immediate reaction and refinement.
Detailed Information Flow Example: Consider Agent 1 in round $t$:
Step 1: Information Gathering
  • Communication Moderator selects peers (e.g., Agent 2 & 3) based on cognitive distance and network topology
  • Agent 1 accesses TMS to understand group consensus and peer specializations
  • Agent 1 receives specific outputs from selected peers
Step 2: Cognitive Processing
  • From TMS: "Agent 2 is performing fact-checking, Agent 3 is identifying mathematical formulas"
  • Agent 1 realizes these tasks are covered, enabling cognitive offloading
  • Agent 1 focuses on alternative solution approaches using Agent 3’s formulas
Step 3: Output Generation and System Update
  • Agent 1 generates new output informed by peer insights and TMS understanding
  • All agent outputs update the TMS for next round
  • Communication Moderator updates network topology based on new cognitive distances
Network Dynamics: The communication network reconstructs dynamically each round, with edges determined by:
  • Fixed in-degree $N$ (CLT working memory constraint)
  • Cognitive similarity for local clustering (reduces extraneous load)
  • Probabilistic rewiring $\beta$ for diverse exploration (distributes intrinsic load)
This dual-channel, dynamic network approach distinguishes CoThinker from existing frameworks that typically use either broadcast communication or static network topologies.
D.3 Prompt Architecture for Agent Parallel Thinking
@@
D.4 Prompt Architecture for Transactive Memory System (TMS) Emulation
As introduced in Section 4.2, CoThinker incorporates a mechanism to emulate a human Transactive Memory System (TMS). A TMS is a collective cognitive resource developed by groups, encompassing a shared understanding of who knows what (metamemory or expertise directory), how to access and integrate this distributed knowledge, and a level of trust in the information provided by different members (Wegner, 1987; Hollingshead, 2001; Lewis, 2003). Effective TMS functioning involves processes of knowledge encoding (assigning information to members or recognizing expertise), storage (individuals retaining specialized knowledge), and retrieval (accessing and using the distributed knowledge), facilitated by member specialization, perceived credibility, and inter-agent coordination (Yoo & Kanawattanachai, 2001). This systematic division and integration of cognitive labor allows groups to handle more complex information and solve problems more effectively than individuals or less coordinated groups.
@@ D.4 Prompt Architecture for Transactive Memory System (TMS) Emulation
D.5 Communication Moderator: Cultivating an Efficient Network via Strong and Weak Ties
The transactional cost induced by inter-agent communication is a key concern in Collaborative CLT (Kirschner et al., 2009; 2018). The Communication Moderator in CoThinker (Section 4.3) strategically structures inter-agent communication by implicitly leveraging principles from social and complex network theories. This design fosters a network optimized for managing cognitive load and enhancing collective intelligence.
@@ D.5 Communication Moderator: Cultivating an Efficient Network via Strong and Weak Ties
For similarity measurement, we employ cosine similarity of text embeddings (e.g., using all-MiniLM-L6-v2) to quantify cognitive proximity between agent outputs.
@@
  • Network Outcome: Introducing such weak ties into a highly clustered network is a hallmark of small-world networks (Watts & Strogatz, 1998). These networks advantageously combine high local clustering with short global average path lengths.
@@
D.6 Synthesizer Module: Consolidation and Cognitive Grounding
The Synthesizer module (Section 4.4) consolidates outputs from all agents ($\{x_{i}^{(T-1)}\}_{i=1}^{M}$) and the final Transactive Memory System state ($\mu^{(T-1)}$) into a single solution for the task $D$ after $T_{max}$ rounds. It can be implemented as an External Agent (dedicated LLM) or an In-group Agent (team member) (Lu et al., 2024; Shinn et al., 2023). The design choice for the Synthesizer can vary, with different cognitive implications drawing from Collaborative Cognitive Load Theory (Kirschner et al., 2018) and Observational Learning (Bandura & Walters, 1977):